OpenStudy (anonymous):

Coffee is draining from a conical filter into a cylindrical coffee pot at the rate of 10 in^3/min. How fast is the level of the pot rising when the coffee in the cone is 5 in. deep. Also, How fast is the level in the cone falling then? Height of cone is 6 inches, diameter of cone is 6 inches and diameter of cylinder is 6 inches too. Please is the answer to the first question 30Pi?

OpenStudy (anonymous):

What um... What subject is this?

OpenStudy (anonymous):

Math @MaydayPaRAYde

OpenStudy (anonymous):

Yes... But... Like Watch TYPE of math?

OpenStudy (anonymous):

It's Calclus..related rates (rate of change)

OpenStudy (anonymous):

Oh Okay! I'll work on it

OpenStudy (anonymous):

Yay!...thank you so much!

OpenStudy (anonymous):

part A is 10/(9 pi). r = 3 for the cone.

OpenStudy (anonymous):

yope!...10/9pi?...how though?

OpenStudy (anonymous):

A.) Diameter of Pot, d = 6 in. Volume = V dV/dt = 10 cu. in./min. Level of Pot = h Find dh/dt: V = π r² h V = π (d² / 4) h V = π [(6)² / 4] h V = π (36 / 4) h V = π (9) h V = 9π h V = 28.27h Differentiating Implicitly Over Time: dV/dt = 28.27(dh/dt) 10 = 28.27(dh/dt) dh/dt = 10 / 28.27 dh/dt = 0.354 The level of the pot is increasing at a rate of 0.354 in./min. (Since the change in volume is the same notwithstanding the sign, the level in the filter makes no difference)

OpenStudy (anonymous):

Hope this was helpful :3 Is there anything else?

OpenStudy (anonymous):

Isn't the volume of a cone 1/3*pi*r^2h

OpenStudy (anonymous):

B.) Diameter of Cone, d = 6 in. Height of Cone, h = 6 in. Volume of Cone = V dVc/dt = - 10 cu. in./min. Find dh/dt: V = 1/3 π r² h V = 1/3 π (d / 2)²h V = 1/3 π (d² / 4)h V = 1/3 π [(6)² / 4)h V = 1/3 π (36 / 4)h V = 2/3 π (9)h V = 18/3 π h V = 6π h V = 18.85h Differentiating: dV/dt = 18.85(dh/dt) - 10 = 18.85(dh/dt) dh/dt = - 10 / 18.85 dh/dt = - 0.531 The level of the cone is decreasing at a rate of 0.531 in./min. Yes ^-^

OpenStudy (anonymous):

hmm...Thank you very much! :-D

OpenStudy (anonymous):

_•) <) )╯ 'where did my other eye go' / \ (•_•) <( (> 'there it is lol' / \

OpenStudy (anonymous):

You're Welcome :DD sorry it took me so long I didn't see your question.

OpenStudy (anonymous):

LOL!!!...

OpenStudy (anonymous):

It's aiit...glad you eventually helped!

OpenStudy (anonymous):

Aha You're welcome :)

OpenStudy (anonymous):

I have found a little problem with your solution @MaydayPaRAYde the question said in part A, that we should find dh/dt when h = 5 inches

OpenStudy (anonymous):

or was it just put to confuse us...just like they gave us the height of the cone? @MaydayPaRAYde

OpenStudy (anonymous):

Maybe I'm just confusing myself